whats (spuare root of x + radical 3) times ( square root of x + 2 radical 3)
Answer by stanbon(75887)

(square root of x + radical 3) times ( square root of x + 2 radical 3)
-----
(sqrt(x)+sqrt(3))(sqrt(x)+2sqrt(3))
Use FOIL:
F: sqrt(x)*sqrt(x) = x
O: sqrt(x)*2sqrt(3) = 2sqrt(3x)
I: sqrt(x)*sqrt(3) = sqrt(3x)
L: sqrt(3)*2sqrt(3) = 2*3 = 6
-----
Ans: x + 3sqrt(3x) + 6
=========================
Cheers,
Stan H.
==========================